consider the following open economy. the real exchange rate is fixed and equal to 1. consumption is given by C=50+0.5(Y-T), investment is I =20, taxes are t=10, and the government budget is balanced. imports and exports are given by I M = 0.2Y and X = 0.02Y* where Y* is foreign GDP.

1
Expert's answer
2021-11-14T17:33:31-0500

(a)

"Y=C+I+G+NX"

"Y=50+0.5(Y-10)+20+10+(0.02Y^*-0.2Y)"

"Y=75+0.5Y+0.02Y^*-0.2Y=75+0.3Y+0.02Y"

"Y=\frac{(75+0.02Y^*)}{0.7}"

(b)

Government purchase multiplier:

"=\frac{1}{1-mpc+mpm}"

"=\frac{1}{(1-0.5+0.2)}=\frac{1}{0.7}=1.43"

If there is no import and export, the the government purchase multiplier:

"=\frac{1}{1-mpc}=\frac{1}{(1-0.5)}=2."

(c)

If "Y^*=10Y"

So,

"Y=\frac{(75+0.02\times 10Y)}{0.7}"

"\implies Y=150"

So, "Y^*=150\times10=1500"

Foreign GDP after adverse demand hit:

"=0.9\times1500=1350"

New domestic GDP:

"=0.1\times1350=135"

So, domestic GDP falls by 10%.

(d)

Government purchase multiplier"=\frac{1}{0.7}=1.43"

Required increase in output = 15.

Required increase in government purchases:

"=\frac{15}{0.7}=21.42"